OpenStudy (anonymous):

simplify. (2r)^4/(2r^4)

zepdrix (zepdrix):

\[\frac{ (2r)^4 }{ 2r^4 }\]This is your problem? :D Hmm, so notice how the 4 is outside of the brackets on the top. It's telling you to raise EVERYTHING inside the brackets to the 4th power. So do that as your first step. :D \[(2r)^4=2^4 *r^4\]